A silo consists of a right circular cylinder surmounted by a right circular 

cone cover. What is the maximum volume of this silo that can be 

constructed using a 100 m2

of plane GI sheet ?


1
Expert's answer
2021-03-25T09:07:57-0400

GIVEN:


Total surface area of given combination=area of given sheet

=S=lateral surface area of cone +curved surface area of cylinder +area of circle


S="100=\pi rl+2\pi rh_2+\pi r^2\\"




as we know,


"l=\sqrt{r^2+h_1^2}"


and "l={r\\over sin\t\\theta}"


therefore,

"S=\pi r\sqrt{r^2+h_2^2}+2\pi rh_2+\pi r^2\\" .......(1)


volume given by,


"V={1\over3}\pi r^2 h_1+\pi r^2h_2" ........(2)



solving 1 and 2 we get,

"h_2=2h_1"

and then ,substituting this in (1),


for volume to be maximum, we diffrentiate V.

"\boxed{{dV\over dr}=0}"


we will get r=4 ,h1= 3 ,h2=6


using the following equation we will get maximum volume as,

350 m3 or say, (351.85)


This is the application of derivatives.